# SynthCoder: A Synthetical Strategy to Tune LLMs for Code Completion

### v2

@@ Abstract
Code completion is a prominent application of Large Language Models (LLMs) in software engineering. Due to the near real-time response requirements of this task, base models with small to medium-sized parameters are typically employed, supplemented by various optimization and post-training techniques. However, these optimization methods often have trade-offs, leading to a seesaw effect where performance improvements on certain datasets or metrics are accompanied by degradations on others – sometimes even falling below the baseline model’s performance. This paper proposes SynthCoder, a model that integrates leading industry practices to achieve state-of-the-art performance on the Fill-in-the-Middle (FIM) code completion task. In specific, we first construct a diverse dataset by combining Abstract Syntax Tree (AST) node extraction with heuristics that simulate developer behavior. Then we enrich our training corpus with cross-file contextual information using the BM25 algorithm and call graphs, enhancing the model’s ability to perform code completion in both file-level and repository-level scenarios. As the last step, we employ a two-stage training process using the Seed-Coder-8B-Base as the base model. First, we fine-tune the model using Curriculum Learning technology. Following this, we perform alignment using Direct Preference Optimization (DPO) with preference pairs generated through Rejection Sampling. Experimental results demonstrate that our final model excels on mainstream repository-level code completion benchmarks, including aiXcoder, ExecRepoBench, CrossCodeEval, and CoLT. Furthermore, our carefully curated training set effectively mitigates the model’s tendency to just repeat existing code, a common issue existing in various code completion models.
@@ I Introduction
Code completion aims to accelerate software development by predicting a developer’s subsequent coding intentions and providing syntax hints and code snippets. Early code completion approaches were primarily based on predefined rules and traditional machine learning methods. They could index the context and offer suggestions as developers typed keywords and identifiers, significantly alleviating the burden of memorizing precise syntax, function names, or Application Programming Interface (API) details [1, 2]. As software systems have grown in complexity, so too have developer expectations for code completion. Traditional methods, however, have struggled to meet these more demanding requirements. In recent years, Large Language Models (LLMs) have brought transformative changes to the field with their powerful capabilities in text understanding, reasoning, and generation [3]. A multitude of products integrating LLMs with code completion have emerged in the market. Examples include the GitHub Copilot [^1] https://github.com/features/copilot plugin, which is compatible with various IDEs, and the Cursor [^2] https://www.cursor.com IDE, which deeply integrates LLMs into various features of the development tool. These products are often built upon proprietary, cloud-based commercial models from providers like OpenAI and Anthropic, and have achieved remarkable success in both performance and reputation. One study demonstrated that using GitHub Copilot can increase programming speed by 55.8% compared to a control group [4]. Furthermore, a survey revealed that developers experience reduced cognitive load and higher job satisfaction after adopting AI-powered coding tools [^3] https://github.blog/news-insights/research/research-quantifying-github-copilots-impact-on-developer-productivity-and-happiness. Concurrently, the industry are also developing open-source alternatives, such as Tabby [^4] https://github.com/TabbyML/tabby.
@@
    We propose a comprehensive methodology for code completion models that organically integrates and improves upon the data construction techniques, and training methods from several existing works [12, 13, 10], resulting in a new state-of-the-art (SOTA) model for FIM code completion.
@@
    We empirically evaluate several mainstream code completion models and discuss the gap between model capabilities and their successful application in completion tools, specifically addressing the issue of models erroneously repeating one or more lines of existing context.
The structure of this paper is as follows. Section III introduces the background and motivation of our research, and Section II reviews related work on applying LLMs to domains such as code completion. Section III provides a detailed account of our entire process, from data acquisition (Section III-A) and data processing (Section III-B) to training details (Section III-C). To validate the effectiveness of our method, in Section IV, we conduct a comprehensive evaluation of all baseline models and SynthCoder on file-level and repository-level code completion tasks, followed by an in-depth analysis of the results. Based on these experimental findings, Section V discusses the gap between code completion performance on test sets versus in practical applications. Section VI addresses the limitations of the current work. Finally, Section VII concludes the paper and suggests directions for future work.
@@ II Related Work
This section describes relevant work from the perspectives of code completion and adopting LLMs for code completion tasks.
@@ II-A Fill in the Middle Code Completion
Large Language Models (LLMs) for code have demonstrated significant potential in code completion tasks. Early works such as StarCoder and CodeLlama laid a foundational basis for this area [7, 6]. However, constructing high-quality training data remains a challenge, particularly in ensuring that the completion target is meaningful in relation to its surrounding context (the prefix and suffix) and covers complex, real-world scenarios. Research in this domain has focused on structuring completion data to enhance model performance in authentic development contexts.
Similarly leveraging code’s structural information, the ExecRepoBench team introduced a multi-level code completion method that also relies on ASTs to build training data. Their approach emphasizes enhancing the model’s capabilities in practical coding scenarios that involve complex, multi-file interdependencies [12]. This reflects a trend in code completion research, which is progressively moving from the single-file level to more complex repository-level scenarios that include cross-file dependencies.
II-B Repository Level Code Completion
@@
II-C Post-training LLMs for Coding
Code generation is a significant application of LLMs. The core technology is founded on the Transformer architecture, which replaced Recurrent Neural Networks (RNNs) [20] and Convolutional Neural Networks (CNNs) [21]. By leveraging a self-attention mechanism, Transformers effectively model long-range dependencies and exhibit excellent scaling stability, laying the groundwork for learning from code corpora [22].
@@ II-C Post-training LLMs for Coding
The decoder-only architecture has become the mainstream choice for code LLMs due to its advantages in text generation, including code. GPT-3 (175B) [23] demonstrated general-purpose capabilities (including code generation) without fine-tuning, and its emergent abilities [24] greatly spurred exploration in the field. Subsequently, OpenAI trained the CodeX model on a vast amount of open-source code (e.g., based on GPT-3.5), revealing for the first time the broad prospects of LLMs in software engineering.
The open-source community has since produced powerful base models, such as Llama/Llama2 [25, 26], the Qwen series, and the DeepSeek series, as well as code-specific models like DeepSeek-Coder [27], Qwen2.5-Coder [8], and Seed-Coder [9]. The base model used for fine-tuning in this paper is from the Seed-Coder series. These models have achieved outstanding results on code benchmarks, with some metrics even surpassing those of larger, closed-source models.
@@
In terms of expanding model capabilities, the o1 model launched by OpenAI in 2024 opened a new dimension: Inference Time Scaling [28]. These models are also known as Reasoning LLMs. Shortly after, the DeepSeek-R1 [29] model was open-sourced, presenting a method for training reasoning models using reinforcement learning that achieved complex task-solving abilities comparable to o1. Although reasoning models are constrained in low-latency tasks like code completion, they are becoming indispensable for more systematic software engineering tasks such as code repair, repository-level code understanding, and complex code generation.
@@
Another critical path to enhancing code LLM performance is the construction of high-quality, domain-specific training data. This is particularly challenging for code completion (the FIM task, as described in Section II-A) and code instruction generation. For the latter, the primary challenge is to increase data diversity and difficulty. Early work like Code Alpaca generated instruction-output pairs from seed tasks using ChatGPT [30], but its data diversity was limited. To address this, subsequent research proposed innovative methods: Evol-Instruct [31] evolutionarily rewrites simple instructions, while OSS-Instruct [32] innovatively uses open-source code as seeds to prompt an LLM to generate relevant problems and solutions.
In summary, current research on code Large Language Models focuses on three key directions: (1) Constructing high-quality Fill in the Middle (FIM) training data by leveraging structural information like Abstract Syntax Trees (ASTs) to enhance the model’s code completion capabilities in both single-file and complex cross-file dependency scenarios [11, 12]; (2) Advancing the code completion task from the single-file level to the Repository Level, which requires an understanding of the overall project structure and cross-file dependencies. Relevant research encompasses innovative paradigms such as retrieval augmentation [15, 16], graph structure modeling [17, 15], and the integration of static analysis [33, 19]; (3) Continuously improving model performance on complex tasks like code generation, understanding, and repair by combining advanced post-training techniques – especially diverse instruction data synthesis methods [31, 32] – with innovative model architectures and optimization strategies [28]. Collectively, these research streams lay the foundation for building more powerful and practical intelligent code assistants. Unfortunately, due to inherent constraints in their internal mechanisms, these different approaches cannot be simply combined to boost the performance of code completion models [34].
@@ III Methodology
The details in training SynthCoder will be elaborated in this section.
@@ III-A Data Retrieval
To build a high-quality and multi-dimensional training corpus for code completion, we constructed data for different application scenarios on two levels. First, we utilized The-Stack-V2 [35] open-source repository dataset to create Fill-in-the-Middle (FIM) code completion samples that include both intra-file and cross-file contexts by masking abstract syntax tree nodes. Second, based on the pre-training corpus from the OpenCoder project, we generated samples for completing function bodies based on function signatures and docstrings. Finally, to improve training efficiency and model performance, we applied a unified quality filtering process to all data, which combined a heuristic rule pipeline with sample selection based on model perplexity.
III-A1 Files from Open-Source Repositories
To construct a code completion dataset with both intra-file and cross-file contexts, with considerations of data quality and the costs of data acquisition, we extracted files in Java, Python, C++, Go, JavaScript, and TypeScript from repositories with 256 or more stars in the the-stack-v2-train-smol-ids [^6] https://huggingface.co/datasets/bigcode/the-stack-v2-train-smol-ids, a subset of The-Stack-V2, which served as the basis for our AST node masking and cross-file context construction.
III-A2 Snippets of Standalone Functions
For the scenario of completing function bodies based on function signatures and docstrings, we used the pre-training annealing data, opc-annealing-corpus [^7] https://huggingface.co/datasets/OpenCoder-LLM/opc-annealing-corpus/, from the OpenCoder project to construct function completion samples. This dataset contains over ten million lines of code snippets synthesized by GPT-4. Specifically, we extracted function implementations from complete code files to serve as the code snippets to be completed, while the function signature, docstring, and other external code related to the target function collectively formed the context. For Java, Python, C++, Go, JavaScript, and TypeScript, we used data from the synthetic_code_snippet subset. As the Go language was not included in this subset, we used the synthetic_qa subset for extraction instead. Since functions in this dataset may lack docstrings, we used the DeepSeek-V3-0324888https://huggingface.co/deepseek-ai/DeepSeek-V3-0324 model to synthesize corresponding docstrings for the extracted functions written in Go.
III-A3 Filtering Based on Data Quality
To improve training efficiency and enhance the final model’s performance, all data from our sources underwent a quality filtering process.
OpenCoder Pipeline
OpenCoder Data Filtering Pipeline is a heuristic filtering framework for large-scale code pre-training corpora, which considers the unique characteristics of different programming languages and includes over 100 filtering rules. The guiding principle of this pipeline is to remove harmful data while ensuring that the overall distribution of the dataset is not significantly affected [13]. In this paper, we adopt this pipeline [^9] https://github.com/OpenCoder-llm/opc_data_filtering to filter the raw code data based on heuristic rules, thereby obtaining high-quality yet diverse code, which in turn improves the overall quality of the subsequently synthesized code completion data.
Perplexity Filtering
@@ III-B Data Construction
The core of our data construction lies in deeply simulating the real-world coding habits and contextual needs of developers. Our method encompasses three key aspects: (1) For constructing the completion target, we combine fine-grained node extraction based on Abstract Syntax Trees (ASTs) with heuristic strategies that mimic user behavior. (2) For building the completion context, we introduce critical cross-file information using code similarity (BM25) and code dependency graphs (Call Graphs). (3) For optimizing data quality, we employ rejection sampling and a multi-stage filtering process to generate high-quality preference pairs for alignment training, aiming to enhance the model’s FIM capabilities and suppress context repetition.
Through this multi-dimensional construction strategy, we aim to generate high-quality training data that more closely resembles real-world application scenarios.
@@ III-B1 Inner File Context
Developers tend to think and code in terms of syntactic nodes that carry specific meaning, rather than handling characters or lines of code in isolation. This observation stems from the top-down, divide-and-conquer strategy prevalent in software development, where developers decompose functionality into multiple, relatively independent code snippets. This approach significantly reduces the cognitive load of developing complex software. After breaking down complex features, developers can concentrate on implementing smaller, logically simpler code segments. This method of progressively decomposing a complex system naturally aligns with how parsing tools break code down into a tree-like structure.
@@
    • Syntax-Aware Triggered Completion: The completion is triggered immediately following a language-specific, syntactically significant token. We define a curated set of such tokens for each programming language, encompassing critical keywords (e.g., for control flow, definitions) and operators (e.g., for assignment, member access). The remainder of the line after the token is designated as the completion target.
  • Code Fragments Within Parentheses (4.86%): Paired parentheses often enclose relatively independent code snippets. In code editors, typing an opening parenthesis automatically generates the closing one and places the cursor in between. We sample data to replicate this common scenario.
  • Code Fragments After Comments (2.92%): When a model cannot directly understand a developer’s intent, the developer may manually write a comment describing the functionality of the code block to be written, then trigger code completion on the next line. To simulate this, we use Tree-sitter to detect full-line comment nodes in code files and treat the first code block immediately following the comment as the completion target.
@@ III-B2 Cross File Context
We use BM25 algorithm [39] to retrieve code snippets similar to the target completion, providing the LLM with inspiration for analogical learning in terms of implementation logic and local code structure. Before computing text similarity, we first segment the code within the project into chunks based on blank lines, ensuring that each snippet is fewer than 20 lines long while remaining relatively self-contained [40]. The query consists of the code to be completed along with some of its adjacent context. We then compute the BM25 similarity score between this query and all code chunks from other files within the project, selecting the highest-scoring snippets (sum up within a maximum total length after tokenization) to serve as cross-file context.
@@
Code import statements directly indicate which external APIs and modules can be called within a file. Retrieving and providing such code information from the project allows the LLM to better decide when to use imported modules and how to use them more correctly.
@@
III-B3 Rejection Sampling
To further optimize the model’s FIM capabilities, this paper introduces paired preference alignment training. The core of this training is to construct a high-quality preference dataset, for which a negative sample generation and filtering pipeline was designed. The main steps are as follows:
Rejection Sampling Strategies
Based on a cross-file code dataset constructed using the BM25 algorithm and call relationship graphs, we use the target model intended for alignment on each sample. With the sampling temperature set to 1.0, we generate 10 candidate negative samples via rejection sampling [10].

### v3

@@ Abstract
Code completion is a prominent application of Large Language Models (LLMs) in software engineering. Due to the near real-time response requirements of this task, base models with small to medium-sized parameters are typically employed, supplemented by various optimization and post-training techniques. However, these optimization methods often have trade-offs, leading to a seesaw effect where performance improvements on certain datasets or metrics are accompanied by degradations on others – sometimes even falling below the baseline model’s performance. This paper proposes SynthCoder, a model that integrates leading industry practices to achieve state-of-the-art performance on the Fill-in-the-Middle (FIM) code completion task. In specific, we first construct a diverse dataset by combining Abstract Syntax Tree (AST) node extraction with heuristics that simulate developer behavior. Then we enrich our training corpus with cross-file contextual information using the BM25 algorithm and call graphs, enhancing the model’s ability to perform code completion in both file-level and repository-level scenarios. As the last step, we employ a two-stage training process using the Seed-Coder-8B-Base as the base model. First, we fine-tune the model using Curriculum Learning technology. Following this, we perform alignment using Direct Preference Optimization (DPO) with preference pairs generated through rejected code sampling. Experimental results demonstrate that our final model excels on mainstream repository-level code completion benchmarks, including aiXcoder, ExecRepoBench, CrossCodeEval, and CoLT. Furthermore, our carefully curated training set effectively mitigates the model’s tendency to just repeat existing code, a common issue existing in various code completion models.
@@ I Introduction
Code completion aims to accelerate software development by predicting a developer’s subsequent coding intentions and providing syntax hints and code snippets. Early code completion approaches were primarily based on predefined rules and traditional machine learning methods. They could index the context and offer suggestions as developers typed keywords and identifiers, significantly alleviating the burden of memorizing precise syntax, function names, or Application Programming Interface (API) details [1, 2]. As software systems have grown in complexity, so too have developer expectations for code completion. Traditional methods, however, have struggled to meet these more demanding requirements. In recent years, Large Language Models (LLMs) have brought transformative changes to the field with their powerful capabilities in text understanding, reasoning, and generation [3]. A multitude of products integrating LLMs with code completion have emerged in the market. Examples include the GitHub Copilot [^1] https://github.com/features/copilot plugin, which is compatible with various IDEs, and the Cursor [^2] https://www.cursor.com IDE, which deeply integrates LLMs into various features of the development tool. These products are often built upon proprietary, cloud-based models from providers like OpenAI and Anthropic, and have achieved remarkable success in both performance and reputation. One study demonstrated that using GitHub Copilot can increase programming speed by 55.8% compared to a control group [4]. Furthermore, a survey revealed that developers experience reduced cognitive load and higher job satisfaction after adopting AI-powered coding tools [^3] https://github.blog/news-insights/research/research-quantifying-github-copilots-impact-on-developer-productivity-and-happiness. Concurrently, the industry are also developing open-source alternatives, such as Tabby [^4] https://github.com/TabbyML/tabby.
@@
    We propose a comprehensive methodology for code completion models that organically integrates and improves upon the data construction and training methods from several existing works [12, 13, 10, 14], resulting in a new state-of-the-art (SOTA) model SynthCoder for FIM code completion based on Seed-Coder-8B-Base.
@@
    We empirically evaluate several mainstream code completion models and discuss the gap between model capabilities and their successful application in completion tools, specifically addressing the issue of models erroneously repeating nearby context.
The structure of this paper is as follows. Section III introduces the background and motivation of our research, and Section II reviews related work on applying LLMs to domains such as code completion. Section III provides a detailed account of our entire process, from data retrieval (Section III-A) and data construction (Section III-B) to details in model post-training(Section III-C). To validate the effectiveness of our method, in Section IV, we conduct a comprehensive evaluation of all baseline models and SynthCoder on file-level and repository-level code completion tasks, followed by an in-depth analysis of the results. Based on these experimental findings, Section V discusses the gap between code completion performance on test sets versus in practical applications. Section VI addresses the limitations of the current work. Finally, Section VII concludes the paper.
@@ II Related Work
This section describes relevant work to position our work.
@@ II-A Fill in the Middle Code Completion
Large Language Models (LLMs) for code have demonstrated significant potential in code completion tasks. The Fill-in-the-Middle (FIM) paradigm, proposed by OpenAI [5], established a specialized input-output format for code completion that has since been widely adopted. Early works such as CodeLlama and StarCoder laid a foundational basis for this area [7, 6]. However, constructing high-quality training data remains a challenge, particularly in ensuring that the completion target is meaningful in relation to its surrounding context and covers complex, real-world scenarios. Research in this domain has focused on structuring completion data to enhance model performance in authentic development contexts [11].
Recent work has further explored syntax-aware FIM pretraining and evaluation. For instance, [15] proposed AST-T5, which incorporates AST structure into pretraining without architectural changes, showing strong performance in code-to-code tasks. [16] introduced SAFIM, a syntax-aware FIM benchmark that evaluates LLMs on structured code completions, revealing that pretraining methods and data quality often outweigh model size. [17] extended this line by proposing structure-aware FIM pretraining with syntax-tree-based masking, though their method focuses on single or sibling AST nodes and lacks open-source implementation. Other studies have addressed tokenization issues in FIM. [18] proposed a character-level infilling method to mitigate sub-token misalignment, which is relevant though not directly focused on syntax trees.
@@
II-B Post-training LLMs for Coding
@@
Code generation is a significant application of LLMs. GPT-3 (175B) [25] demonstrated general-purpose capabilities (including code generation) without fine-tuning, and its emergent abilities [26] greatly spurred exploration in the field. Subsequently, OpenAI trained the CodeX model on a vast amount of open-source code (e.g., based on GPT-3.5), revealing for the first time the broad prospects of LLMs in software engineering. The open-source community has since produced powerful base models specifically for code-related tasks, like the DeepSeek-Coder [27], Qwen2.5-Coder [8], and Seed-Coder [9] series. These models have achieved outstanding results on code benchmarks, with some metrics even surpassing those of larger, closed-source models.
@@
In terms of expanding model capabilities, the o1 model launched by OpenAI in 2024 opened a new dimension: Inference Time Scaling [28]. These models are also known as Reasoning LLMs. Shortly after, the DeepSeek-R1 [29] model was open-sourced, presenting a method for training reasoning models using reinforcement learning that achieved complex task-solving abilities comparable to o1. Although reasoning models are constrained in low-latency tasks like code completion, they are becoming indispensable for more systematic software engineering tasks such as code repair, repository-level code understanding, and complex code generation. While the field of code generation is evolving towards autonomous agents—benchmarked by SWE-bench [30] and embodied by frameworks like SWE-agent [31] and OpenDevin [32]—their iterative, tool-heavy nature induces significant latency. This makes them unsuitable for real-time code completion.
@@
Another critical path to enhancing code LLM performance is the construction of high-quality, domain-specific training data. This is particularly challenging for code completion (the FIM task, as described in Section II-A) and code instruction generation. For the latter, the primary challenge is to increase data diversity and difficulty. Early work like Code Alpaca generated instruction-output pairs from seed tasks using ChatGPT [33], but its data diversity was limited. To address this, subsequent research proposed innovative methods: Evol-Instruct [34] evolutionarily rewrites simple instructions, while OSS-Instruct [35] innovatively uses open-source code as seeds to prompt an LLM to generate relevant problems and solutions. Meanwhile, [36] aligned FIM training with code generation tasks, demonstrating the versatility of FIM-based training paradigms beyond completion.
@@ III Methodology
The details in the methodology SynthCoder will be elaborated in this section.
@@ III-A Data Retrieval
We started with the an officially deduplicated subset [^6] https://huggingface.co/datasets/bigcode/the-stack-v2-train-smol-ids of Stack v2 [37] dataset, which is one of the biggest code corpora. To balance data quality and acquisition cost, we curated a specialized subset by selecting files from repositories with 256 or more stars, focusing on six languages: Java (25%), Python (25%), Go (14%), JavaScript (14%), C++ (14%), and TypeScript (8%). This final collection served as the basis for our Fill-in-the-Middle (FIM) training set construction. The raw data was subsequently filtered through a two-stage filtering process: a heuristic rule-based pipeline and a model perplexity-based sample selection.
III-A1 Filtering Based on Heuristics
In this work, we also OpenCoder’s opc_data_filtering [^7] https://github.com/OpenCoder-llm/opc_data_filtering, which is a heuristic filtering framework for large-scale code pre-training corpora, which considers the unique characteristics of different programming languages and includes over 100 filtering rules [13], with guiding principle to remove harmful data while ensuring that the overall distribution of the dataset is not significantly affected, thus obtaining a code dataset with relatively high quality yet remain diversity.
III-A2 Filtering Based on Perplexity
@@ III-B Data Construction
The core objective of our dataset construction is to simulate, to the greatest extent possible, the coding practices, contextual needs, and code completion expectations of human engineers in real-world file-level and repository-level programming tasks. This section elaborates the process to construct such a training corpora to meet this objective.
@@ III-B1 Inner File Context
Developers tend to think and code in terms of syntactic nodes that carry specific meaning, rather than handling characters or lines of code in isolation. This observation stems from the top-down, divide-and-conquer strategy prevalent in software development, where developers decompose functionality into multiple, relatively independent code snippets. This approach significantly reduces the cognitive load of developing complex software. After breaking down complex features, developers can concentrate on implementing smaller, logically simpler code segments.
@@
    • Syntax-Aware Triggered Completion: The completion is triggered immediately following a language-specific, syntactically significant token. We define a curated set of such tokens for each programming language, encompassing critical keywords (e.g., for control flow, definitions) and operators (e.g., for assignment, member access), which the remainder of the line after these token will be designated as the completion target.
  • Code Fragments Within Parentheses (4.86%): In code editors, typing an opening parenthesis automatically generates the closing one and places the cursor in between. We sample data to replicate this common scenario.
  • Code Fragments After Comments (2.92%): The developer may manually write a comment describing the code block to be written, then trigger completion on the next line. To simulate this, we detect full-line comment nodes in code files and treat the first code block immediately following the comment as the completion target.
@@ III-B2 Cross File Context
We use BM25 algorithm [41] to retrieve code snippets similar to the target completion, providing the LLM with inspiration for analogical learning in terms of implementation logic and local code structure. Before computing text similarity, we first segment the code within the project into chunks based on blank lines, ensuring that each snippet is fewer than 20 lines long while remaining relatively self-contained [42]. The query consists of the code to be completed along with some of its adjacent context. We then compute the BM25 similarity score between this query and all code chunks from other files within the project, selecting the highest-scoring snippets (sum up within a maximum total length after tokenization) to serve as the retrieval based-part of cross-file context.
@@
Code import statements directly indicate which external APIs and modules can be called within a file. Retrieving and providing such code information from the project allows the LLM to better decide when to and how to use the imported snippets correctly.
@@
III-B3 Rejected Code Sampling
To further optimize the model’s FIM capabilities, this paper introduces paired preference alignment training. The core of this training is to construct a high-quality preference dataset, for which a negative sample generation and filtering pipeline was designed.
The main steps are as follows:
Rejected Code Sampling Strategies
Based on the dataset containing cross-file information, we use the best checkpoint in the finetuning stage which is the model intended for alignment to generate 10 candidate negative samples via rejected code sampling [10] on each sample with temperature set to 1.0.